OpenStudy (anonymous):

someone please help :( 9x - 7 = -7

OpenStudy (anonymous):

you have to solve it its like a algebraic equation. im just really stuck on it

OpenStudy (♪chibiterasu):

Just add 7 on both sides.

OpenStudy (koikkara):

\(9x - 7 = -7\), x= 0/9 =infinity

OpenStudy (anonymous):

i did but then i got stuck

OpenStudy (♪chibiterasu):

After that, you divide both sides by nine.

OpenStudy (anonymous):

thanks guys :)

OpenStudy (♪chibiterasu):

0/9 is just 0

OpenStudy (koikkara):

But its wrong...yeah, its...zero lol

OpenStudy (anonymous):

thats what i had trouble with. thank you so much